OpenStudy (anonymous):

Some help me please, will medal and fan?! What is the equation of a line that passes through (4, 7) and has a slope of −14?

rishavraj (rishavraj):

\[\frac{ y - y _{1} }{ x - x _{1} } = m\]

OpenStudy (anonymous):

A)\[y=\frac{ x }{ 4 }-\frac{ 7 }{ 4 }\] B) \[y=-\frac{ x }{ 4 }+\frac{ 7 }{ 4 }\] C) \[y=-\frac{ x }{ 4 }+8\] D) \[y=\frac{ x }{ 4 }+8\]

OpenStudy (anonymous):

what is the slope? (-4)

OpenStudy (anonymous):

the equation is y =m(x)+c put in slope m,and x given =4 and y given =7 and find c

OpenStudy (anonymous):

I got \[C=-\frac{ 7 }{ 16 }\] @jamaica25 @rishavraj

rishavraj (rishavraj):

@itsdaniela whts the slope ???? is it -4 ????

OpenStudy (anonymous):

yes thats what i used @rishavraj

rishavraj (rishavraj):

then 7 = 4 (-4) + C 7 = -16 + C C = 23

rishavraj (rishavraj):

According to ur question

OpenStudy (anonymous):

So which one is right? @rishavraj

rishavraj (rishavraj):

i m confused ....r u sure tht the question or its options r correctly mentioned?????

OpenStudy (phi):

it looks like the given slope is -1/4 Not -14

rishavraj (rishavraj):

@phi awesome...:))) its really supposed to be \[-\frac{ 1 }{ 4 }\] @itsdaniela use slope = -1/4 ..and solve it..

OpenStudy (anonymous):

Using the same equation? @rishavraj

rishavraj (rishavraj):

yup u can use any one of the two \[\frac{ y -y_1 }{ x - x_1 } = m\] or y = mx + c

OpenStudy (anonymous):

Thank you soo much! @rishavraj

rishavraj (rishavraj):

u r most welcome,....:)